What is the concentration of cells in the starting sample if 1 ml was plated per dilution and the starting sample was river water?

Report your answer below as a decimal with one significant digit after the decimal and report the exponent the separate box. See this example, if your answer is 5.2 x 108 you would report it as "5.2" You have the following results from a serial dilution experiment: Sample Dilution Plate Count 1:101 Too many to count 1:102 T
Answer #1

CFU/mL= number of colonies/ ( volume plated in mL\times dilution)

dilution= 1/10^-4

number of colonies in the range of 30- 300 is used for the CFU/ mL calculation, here for dilution 1:10^4 the number of colonies=287

volume plated=1 mL

CFU/mL= 287/(1\times10^-4)

= 2.87\times10^6 /mL

concentration= 2.87\times10^6 cells/mL